# Democratizing Text-to-Image Masked Generative Models with Compact Text-Aware One-Dimensional Tokens

### v1

@@ 1 Introduction
In recent years, text-to-image generation has seen remarkable progress across various frameworks, including diffusion models [50, 51, 23], autoregressive visual models [26, 58, 57], and masked generative models [12, 7, 24]. A crucial component of these models is a robust image tokenizer—either discrete [61] or continuous [35]—which transforms images into tokenized representations. These models then incorporate text conditions with the tokenized representations using methods such as cross-attention [50], concatenation [8], or conditioning embeddings [44], ultimately leveraging the tokenizer to generate high-quality images aligned with input text prompts.
@@
In this work, we address this question by building on the recent concept of compact one-dimensional tokenizers to facilitate both efficient and high-quality image tokenization. Traditional tokenizers [35, 61, 22] rely on 2D grid-based latent representations, which struggle to handle the inherent redundancies in images, as neighboring patches often display similarities. Recently, Yu et al. [69] introduced the Transformer-based 1-Dimensional Tokenizer (TiTok), which efficiently tokenizes images into compact 1D latent sequences by removing the fixed correspondence between latent representation and 2D image patches (i.e., each 1D token can represent any region in an image, rather than being tied to a specific patch). This approach results in a highly efficient tokenizer, significantly improving sampling speed compared to previous methods [61, 22, 11, 44].
@@
Notably, despite being trained entirely on publicly available datasets, MaskGen achieves strong performance and efficiency in text-to-image generation. On MJHQ-30K [36], the lighter MaskGen-L (568M) with discrete tokens achieves a generation FID of 7.74, outperforming Show-o [63] (14.99) with 30.3× faster inference throughput. It also surpasses SD-2.1 [50] (26.96) and PixArt-$\alpha$ [14] (9.85) while requiring only 2% and 21% of their training times, respectively. Furthermore, the larger MaskGen-XL (1.1B) achieves FID scores of 7.51 and 6.53 on MJHQ-30K [36] and overall scores of 0.57 and 0.55 on the GenEval [27] benchmark using discrete and continuous tokens, respectively.
@@ 3 Preliminary
TiTok [69] is a transformer-based, 1-dimensional Vector-Quantized (VQ) tokenizer that diverges from traditional 2D grid-based latent space tokenization, instead opting for a compact 1D representation that bypasses 2D spatial structure preservation. Given an input image $\mathbf{I}\in\mathbb{R}^{H\times W\times 3}$, the tokenization phase of TiTok involves downscaling the image by a factor of $f$, resulting in patches $\mathbf{P}\in\mathbb{R}^{\frac{H}{f}\times\frac{W}{f}\times D}$. These patches are concatenated with a set of latent tokens $\mathbf{L}\in\mathbb{R}^{K\times D}$. The combined sequence is then passed through a Vision Transformer (ViT) [20] encoder, $Enc$, to generate embeddings. Only the embeddings corresponding to the latent tokens, $\mathbf{Z}_{1D}\in\mathbb{R}^{K\times D}$, are retained, forming a compact 1D latent representation. This representation is then quantized through a quantizer $Quant$ by mapping it to the nearest codes in a learnable codebook.
In the de-tokenization phase, TiTok uses a sequence of mask tokens $\mathbf{M}\in\mathbb{R}^{\frac{H}{f}\times\frac{W}{f}\times D}$, which are concatenated with the quantized codes. The resulting sequence is processed by a ViT decoder, $Dec$, to reconstruct the image $\mathbf{\hat{I}}$. Formally, with $\oplus$ denoting concatenation, the tokenization and de-tokenization processes in TiTok can be represented as:
$\mathbf{Z}_{1D}=Enc(\mathbf{P}\oplus\mathbf{L}),$
$\mathbf{\hat{I}}=Dec(Quant(\mathbf{Z}_{1D})\oplus\mathbf{M}).$
@@
Masked Generative Models with Continuous Tokens maintain a conceptual similarity to discrete-token models but operate on continuous tokens, which reduces information loss from quantization. Recently, the diffusion loss [39, 24] was introduced, enabling these models to approximate the distribution of each image token independently. In this framework, Transformers generate a conditioning vector for each masked token, which is then input to a small multi-layer perceptron (MLP) that learns a denoising function [33] conditioned on it. This per-token conditioning and denoising allow the sampling process to directly apply to the probability distribution of each token [39].
@@ 4.1 TA-TiTok: Text-aware 1D Tokenizer
Extending TiTok to Support Continuous Tokens. To maximize the efficiency of the TiTok framework in diffusion-based models, we extend its discrete 1D tokens to continuous 1D VAE representations. Rather than using a quantizer $Quant$ to map the 1D latents $\mathbf{Z}_{1D}$ to the nearest codebook entries, we model $\mathbf{Z}_{1D}$ as a Gaussian distribution and apply KL divergence regularization, resulting in a compact 1D VAE representation. This continuous representation retains the efficiency and structure of the TiTok framework, consistently improving reconstruction quality by avoiding the information loss associated with quantization. Moreover, this KL variant of TiTok integrates seamlessly with diffusion models, serving as a drop-in replacement for standard 2D VAEs [50, 55, 39]. In the supplementary materials (Tab. 14 in Appendix Sec. F), we validate this approach using the state-of-the-art image generation model MAR [39] on ImageNet [17], where our modification achieves a significant reduction in training costs and an increase in inference speed, all while maintaining comparable performance. This enhancement thus contributes to both the efficiency and flexibility of diffusion-based generation. For clarity, we refer to both variants collectively as TiTok, using the labels VQ and KL to denote the discrete and continuous versions, respectively, in the following text.
@@
As shown in Fig. 2(a), the tokenization stage in TA-TiTok mirrors that of TiTok, transforming images into compact 1D tokens, either discrete or continuous. In the de-tokenization stage, however, TA-TiTok incorporates text guidance by using text embeddings generated by a pre-trained vision-language model (e.g., CLIP’s text encoder [47]). These text embeddings are projected through a linear layer to align with the channel dimensions of our TA-TiTok’s ViT decoder, resulting in $\mathbf{T}\in\mathbb{R}^{N\times D}$, where $N$ is the number of context tokens predefined by the vision-language model (e.g., 77 for CLIP’s text encoder). This text embedding is then concatenated with the latent tokens $\mathbf{Z}_{1D}$ and the mask tokens $\mathbf{M}$ before passing through the decoder $Dec$, yielding the reconstructed image $\mathbf{\hat{I}}$.
Formally, the de-tokenization phase of the VQ variant can be expressed as:
$\mathbf{\hat{I}}=Dec(Quant(\mathbf{Z}_{1D})\oplus\mathbf{T}\oplus\mathbf{M}).$
For the KL variant, the de-tokenization phase follows a similar formulation but omits the quantizer $Quant$, as it operates on continuous representations directly.
@@
The design of TA-TiTok incorporates text tokens exclusively into the tokenizer decoder with minimal modifications. To validate this approach, we also experimented with adding text tokens to both the encoder and decoder. The results showed similar performance to the decoder-only approach, suggesting that incorporating textual information during de-tokenization is sufficient for capturing high-level semantic information. Based on these findings, we adopt the simpler decoder-only design. For additional details, see Tab. 9 in Appendix Sec. E.
Figure 3: Overview of MaskGen. MaskGen is a family of text-to-image masked generative models that supports both discrete (VQ variant) and continuous (KL variant) token representations. For discrete tokens, MaskGen is trained with cross-entropy loss [11], while for continuous tokens, it employs diffusion loss [39]. The architecture is designed by concatenating text conditions with TA-TiTok’s latent tokens (both masked and unmasked) and feeding them into Diffusion Transformer blocks [44], with separate adaptive LayerNorm (adaLN) layers for text and image modalities, following MM-DiT [23]. Additionally, aesthetic scores are incorporated as conditioning signals via adaLN. To encode captions, MaskGen uses the CLIP text encoder [47] instead of the more resource-intensive T5-XXL [48], making it more accessible to research groups with limited computational resources.
@@ 4.2 MaskGen: T2I Masked Generative Model
To fully leverage TA-TiTok’s capabilities, we propose MaskGen, a family of text-to-image (T2I) masked generative models. As illustrated in Fig. 3, MaskGen utilizes TA-TiTok to tokenize images into tokens and a CLIP [47] text encoder to extract both global and pooled text embeddings. Inspired by [23], we concatenate the global text embedding with the image tokens, feeding this combined sequence into multi-modal Diffusion Transformer (MM-DiT) blocks [23, 44] for attention operations. To accommodate the distinct properties of text and image embeddings, we apply separate adaptive LayerNorm (adaLN [6, 44]) layers, where the scale and shift parameters are computed based on the sum of the embedding vectors of the pooled text embedding and the time step embedding. We note that while a more powerful text encoder, such as T5-XXL [48], could be directly integrated into the pipeline to enhance performance, as demonstrated in studies like [51, 12, 23], we choose CLIP for its computational efficiency and reduced storage demands, making MaskGen more accessible in resource-constrained settings.
We also incorporate aesthetic scores [53] as another condition by projecting them similarly into sinusoidal embeddings and appending them to the pooled text embedding. This feature provides an extra layer of control over the image quality and style during sampling, making the generation process more adaptable and flexible.
@@ 5 Experimental Results
In this section, we provide the implementation details of TA-TiTok and MaskGen in Sec. 5.1. We then present ablation studies on key design choices in Sec. 5.2. The main results are further discussed in Sec. 5.3.
@@ 5.1 Implementation Details
We implement TA-TiTok based on TiTok [69] with minimal modifications. We then introduce MaskGen model variants and the datasets used for training, along with the training and evaluation protocols applied in our study.
TA-TiTok Model Variants. We present three variants of our TA-TiTok, each containing $K$ = 32, 64, or 128 1D latent tokens, following the architecture of TiTok. Both the tokenizer and de-tokenizer utilize a patch size of $f=16$. For the VQ variant, the codebook is configured with 8192 entries, where each entry is a vector with 64 channels. For the KL variant, we use a continuous embedding with 16 channels, following MAR [39]. For the encoder ($Enc$) and decoder ($Dec$), we find that increasing the size of $Dec$ to ViT-L [20] is beneficial when training on large-scale datasets across all variants. For $Enc$, we adopt ViT-B for all variants except the KL variant with 128 tokens, where ViT-S is sufficient.
@@
MaskGen Model Variants. We introduce two variants of MaskGen: MaskGen-L (large, 568M parameters) and MaskGen-XL (extra-large, 1.1B parameters), with configurations detailed in Tab. 1. For continuous token processing in MaskGen, we incorporate an additional DiffLoss MLP [39], comprising 8 MLP layers with channel sizes aligned to the transformer’s, adding an extra 44M and 69M parameters for MaskGen-L and MaskGen-XL, respectively. To offset the additional training and inference cost introduced by the DiffLoss MLP, we use 128 tokens for the discrete MaskGen variant and 32 tokens for the continuous MaskGen variant. Ablation studies on the impact of token count for MaskGen are provided in the Appendix.
Dataset (recaption). For training TA-TiTok and MaskGen, we utilized various open-source datasets: DataComp-1B [25], CC12M [13], LAION-aesthetic [1], LAION-art [3], LAION-pop [4], JourneyDB [56], and DALLE3-1M [21]. All training images are filtered to ensure their longer side is greater than 256 pixels and their aspect ratio is less than 2. TA-TiTok is trained exclusively with DataComp-1B. MaskGen undergoes a two-stage training process: first, it is pre-trained for image-text alignment using DataComp-1B, CC12M, and LAION-aesthetic, where DataComp-1B and CC12M are additionally filtered to include only images with aesthetic scores [2] higher than 5.0. Following pre-training, MaskGen is fine-tuned using aesthetic images from LAION-art, LAION-pop, JourneyDB, DALLE3-1M, and DataComp-1B (filtered for aesthetic scores above 6.0). For LAION-art, LAION-pop, and the filtered DataComp-1B datasets, we enhance the image captions using Molmo [16], employing one of four different prompts along with the original caption. Details regarding the dataset preparation and recaptioning process are provided in the Appendix.
Training. We adhere strictly to the hyperparameters used to train TiTok across all TA-TiTok variants. Specifically, TA-TiTok is trained with a batch size of 1024 for 1 epoch (650k steps) on the filtered DataComp dataset, using a maximum learning rate of 1e-4 and a cosine learning rate schedule. For MaskGen with discrete tokens, we employ a batch size of 4096, leveraging weight tying [46] to stabilize training, with a cosine learning rate schedule and a maximum learning rate of 4 × 10-4. For MaskGen with continuous tokens, to accommodate the diffusion loss, we use a constant learning rate schedule with a maximum rate of 1 × 10-4 and a batch size of 2048. Masked tokens are sampled by randomly selecting the masking rate from [0, 1] on a cosine schedule, following MaskGIT [11], and text conditioning is randomly dropped with a 0.1 probability to enable classifier-free guidance [32]. Full training hyperparameters are in the Appendix.
Evaluation. Our evaluation pipeline closely follows prior works [65, 12]. The images are generated from text prompts without rejection sampling, and classifier-free guidance [32] has been used to enhance generation quality. Unless specified otherwise, MaskGen uses 16 and 32 sampling steps for VQ and KL architectures, respectively. To assess different aspects of our models’ performance, we employ multiple evaluation metrics. For TA-TiTok, we measure reconstruction quality using reconstruction FID (rFID) [31] and inception score (IS) [52] on ImageNet [17] validation set. For MaskGen’s text-to-image generation capabilities, we utilize a comprehensive set of metrics: FID on MJHQ [37] to assess aesthetic quality, and GenEval [27] score to measure the alignment between text prompts and their corresponding generated images.
5.2 Ablation Studies
@@
Improved One-Stage Training Recipe. Tab. 2 summarizes the performance gains achieved by our improved one-stage training recipe compared to the original training schemes in [69]. As shown, the adopted one-stage training significantly outperforms the original one-stage training with an rFID improvement of 2.72.

### v2

@@ Abstract
Equal contribution. Work done as a research intern at ByteDance.
@@ 1 Introduction
In recent years, text-to-image generation has seen remarkable progress across various frameworks, including diffusion models [54, 55, 23, 58, 53, 30, 69], autoregressive visual models [26, 63, 62, 52, 51], and masked generative models [12, 7, 24]. A crucial component of these models is a robust image tokenizer—either discrete [66] or continuous [36]—which transforms images into tokenized representations. These models then incorporate text conditions with the tokenized representations using methods such as cross-attention [54], concatenation [8], or conditioning embeddings [45], ultimately leveraging the tokenizer to generate high-quality images aligned with input text prompts.
@@
In this work, we address this question by building on the recent concept of compact one-dimensional tokenizers to facilitate both efficient and high-quality image tokenization. Traditional tokenizers [36, 66, 22] rely on 2D grid-based latent representations, which struggle to handle the inherent redundancies in images, as neighboring patches often display similarities. Recently, Yu et al[74] introduced the Transformer-based 1-Dimensional Tokenizer (TiTok), which efficiently tokenizes images into compact 1D latent sequences by removing the fixed correspondence between latent representation and 2D image patches (i.e., each 1D token can represent any region in an image, rather than being tied to a specific patch). This approach results in a highly efficient tokenizer, significantly improving sampling speed compared to previous methods [66, 22, 11, 45].
@@
[…]
@@ 3 Preliminary
TiTok [74] is a transformer-based, 1-dimensional Vector-Quantized (VQ) tokenizer that diverges from traditional 2D grid-based latent space tokenization, instead opting for a compact 1D representation that bypasses 2D spatial structure preservation. Given an input image $\mathbf{I}\in\mathbb{R}^{H\times W\times 3}$, the tokenization phase of TiTok involves downscaling the image by a factor of $f$, resulting in patches $\mathbf{P}\in\mathbb{R}^{\frac{H}{f}\times\frac{W}{f}\times D}$. These patches are concatenated with a set of latent tokens $\mathbf{L}\in\mathbb{R}^{K\times D}$. The combined sequence is then passed through a Vision Transformer (ViT) [20] encoder, $\mathtt{Enc}$, to generate embeddings. Only the embeddings corresponding to the latent tokens, $\mathbf{Z}_{\text{1D}}\in\mathbb{R}^{K\times D}$, are retained, forming a compact 1D latent representation. This representation is then quantized through a quantizer $\mathtt{Quant}$ by mapping it to the nearest codes in a learnable codebook.
In the de-tokenization phase, TiTok uses a sequence of mask tokens $\mathbf{M}\in\mathbb{R}^{\frac{H}{f}\times\frac{W}{f}\times D}$, which are concatenated with the quantized codes. The resulting sequence is processed by a ViT decoder, $\mathtt{Dec}$, to reconstruct the image $\mathbf{\hat{I}}$. Formally, the tokenization and de-tokenization processes in TiTok can be represented as:
[_$[\_~;\mathbf{Z}_{\text{1D}}]=\mathtt{Enc}([\mathbf{P};\mathbf{L}]),$ (1)
[_$[\_~;\mathbf{\hat{I}}]=\mathtt{Dec}([\mathtt{Quant}(\mathbf{Z}_{\text{1D}});\mathbf{M}]),$ (2)
where [;][;][ ; ] denotes concatenation along the sequence dimension, and _\__ represents ignored tokens that are not used in subsequent operations.
@@
Masked Generative Models with Continuous Tokens maintain a conceptual similarity to discrete-token models but operate on continuous tokens, which reduces information loss from quantization. Recently, the diffusion loss [40, 24] was introduced, enabling these models to approximate the distribution of each image token independently. In this framework, Transformers generate a conditioning vector for each masked token, which is then input to a small multi-layer perceptron (MLP) that learns a denoising function [34] conditioned on it. This per-token conditioning and denoising allow the sampling process to directly apply to the probability distribution of each token [40]. Unlike discrete-token models, predicted tokens are masked randomly in each iteration.
@@ 4.1 TA-TiTok: Text-aware 1D Tokenizer
Extending TiTok to Support Continuous Tokens. To maximize the efficiency of the TiTok framework in diffusion-based models, we extend its discrete 1D tokens to continuous 1D VAE representations. Rather than using a quantizer $\mathtt{Quant}$ to map the 1D latents $\mathbf{Z}_{\text{1D}}$ to the nearest codebook entries, we model $\mathbf{Z}_{\text{1D}}$ as a Gaussian distribution and apply KL divergence regularization, resulting in a compact 1D VAE representation. This continuous representation retains the efficiency and structure of the TiTok framework, consistently improving reconstruction quality by avoiding the information loss associated with quantization. Moreover, this KL variant of TiTok integrates seamlessly with diffusion models, serving as a drop-in replacement for standard 2D VAEs [54, 60, 40]. In the supplementary materials (Tab. 13 in the Appendix), we validate this approach using the state-of-the-art image generation model MAR [40] on ImageNet [17], where our modification achieves a significant reduction in training costs and an increase in inference speed, all while maintaining comparable performance. This enhancement thus contributes to both the efficiency and flexibility of diffusion-based generation. For clarity, we refer to both variants collectively as TiTok, using the labels VQ and KL to denote the discrete and continuous versions, respectively, in the following text.
@@
As shown in Fig. 2(a), the tokenization stage in TA-TiTok mirrors that of TiTok, transforming images into compact 1D tokens, either discrete or continuous. In the de-tokenization stage, however, TA-TiTok incorporates text guidance by using text embeddings generated by a pre-trained vision-language model (e.g., CLIP’s text encoder [48]). These text embeddings are projected through a linear layer to align with the channel dimensions of our TA-TiTok’s ViT decoder, resulting in $\mathbf{T}\in\mathbb{R}^{N\times D}$, where $N$ is the number of context tokens predefined by the vision-language model (e.g., 77 for CLIP’s text encoder). This text embedding is then concatenated with the latent tokens $\mathbf{Z}_{\text{1D}}$ and the mask tokens $\mathbf{M}$ before passing through the decoder $\mathtt{Dec}$, yielding the reconstructed image $\mathbf{\hat{I}}$.
Formally, the de-tokenization phase of the VQ variant of TA-TiTok can be expressed as:
[_;_$[\_~;\_~;\mathbf{\hat{I}}]=\mathtt{Dec}([\mathtt{Quant}(\mathbf{Z}_{\text{1D}});\mathbf{T};\mathbf{M}]).$ (3)
For the KL variant, the de-tokenization phase follows a similar formulation but omits the quantizer $\mathtt{Quant}$, as it operates on continuous representations directly.
@@
The design of TA-TiTok incorporates text tokens exclusively into the tokenizer decoder with minimal modifications. To validate this approach, we also experimented with adding text tokens to both the encoder and decoder. The results showed similar performance to the decoder-only approach, suggesting that incorporating textual information during de-tokenization is sufficient for capturing high-level semantic information. Based on these findings, we adopt the simpler decoder-only design. For additional details, see Tab.10 in the Appendix.
Figure 3: Overview of MaskGen. MaskGen is a family of text-to-image masked generative models that supports both discrete (VQ variant) and continuous (KL variant) token representations. For discrete tokens, MaskGen is trained with cross-entropy loss [11], while for continuous tokens, it employs diffusion loss [40]. The architecture is designed by concatenating text conditions with TA-TiTok’s latent tokens (both masked and unmasked) and feeding them into Diffusion Transformer blocks [45], with separate adaptive LayerNorms (adaLN), linear projections, and feedforward networks (FFN) for text and image modalities, following MM-DiT [23]. Additionally, aesthetic scores are incorporated as conditioning signals via adaLN. To encode captions, MaskGen uses the CLIP text encoder [48] instead of the more resource-intensive T5-XXL [49], making it more accessible to research groups with limited computational resources.
@@ 4.2 MaskGen: T2I Masked Generative Model
To fully leverage TA-TiTok’s capabilities, we propose MaskGen, a family of text-to-image (T2I) masked generative models. As illustrated in Fig. 3, MaskGen utilizes TA-TiTok to tokenize images into tokens and a CLIP [48] text encoder to extract both global and pooled text embeddings. Inspired by [23], we concatenate the global text embedding with the image tokens, feeding this combined sequence into multi-modal Diffusion Transformer (MM-DiT) blocks [23, 45] for attention operations. To accommodate the distinct properties of text and image embeddings, we apply separate adaptive LayerNorm (adaLN [6, 45]) layers, where the scale and shift parameters are computed based on the pooled text embedding. We note that while a more powerful text encoder, such as T5-XXL [49], could be directly integrated into the pipeline to enhance performance, as demonstrated in studies like [55, 12, 23], we choose CLIP for its computational efficiency and reduced storage demands, making MaskGen more accessible in resource-constrained settings.
We also incorporate aesthetic scores [57] as another condition by projecting them into sinusoidal embeddings and appending them to the pooled text embedding. This feature provides an extra layer of control over the image quality and style during sampling, making the generation process more adaptable and flexible.
@@ 5.1 Implementation Details
TA-TiTok Model Variants. We present three variants of our TA-TiTok, each containing $K$ = 32, 64, or 128 1D latent tokens, following the architecture of TiTok. Both the tokenizer and de-tokenizer utilize a patch size of $f=16$. For the VQ variant, the codebook is configured with 8192 entries, where each entry is a vector with 64 channels. For the KL variant, we use a continuous embedding with 16 channels, following MAR [40]. For the encoder ($\mathtt{Enc}$) and decoder ($\mathtt{Dec}$), we find that increasing the size of $\mathtt{Dec}$ to ViT-L [20] is beneficial when training on large-scale datasets across all variants. For $\mathtt{Enc}$, we adopt ViT-B for all variants except the KL variant with 128 tokens, where ViT-S is sufficient.
@@
MaskGen Model Variants. We introduce two variants of MaskGen: MaskGen-L (568M parameters) and MaskGen-XL (1.1B parameters), with configurations detailed in Tab. 1. For continuous token processing in MaskGen, we incorporate an additional DiffLoss MLP [40], comprising 8 MLP layers with channel sizes aligned to the transformer’s, adding an extra 44M and 69M parameters for MaskGen-L and MaskGen-XL, respectively. To offset the additional training and inference cost introduced by the DiffLoss MLP, we use 128 tokens for the discrete MaskGen variant and 32 tokens for the continuous MaskGen variant.
Training Data. For training TA-TiTok and MaskGen, we utilized various open-source datasets: DataComp-1B [25], CC12M [13], LAION-aesthetic [1], LAION-art [3], LAION-pop [4], JourneyDB [61], and DALLE3-1M [21]. TA-TiTok is trained exclusively with DataComp-1B. MaskGen undergoes a two-stage training process: pre-training for image-text alignment and fine-tuning using aesthetic images. Details regarding the dataset preparation and recaptioning process are provided in the Appendix.
Evaluation Metrics. Our evaluation pipeline closely follows prior works [71, 12]. The images are generated without rejection sampling, and classifier-free guidance [33] is used to enhance generation quality. Unless specified otherwise, MaskGen uses 16 and 32 sampling steps for VQ and KL variants, respectively. For TA-TiTok, we measure reconstruction quality using reconstruction FID (rFID) [32] and inception score (IS) [56] on ImageNet [17] validation set. For MaskGen’s text-to-image generation capabilities, we utilize FID on MJHQ [38] to assess aesthetic quality, and GenEval [27] score to measure the alignment between text prompts and their corresponding generated images.
@@
5.2 Optimized Image Tokenization with TA-TiTok
Improved One-Stage Training Recipe. Tab. 2 summarizes the performance gains of our improved one-stage training recipe over the original schemes in [74]. As observed, the adopted one-stage training significantly outperforms the original, achieving an rFID improvement of 2.72.
Figure 4: Visualization of Latent Token Attention Map and Latent Code Swapping. The results are from VQ variant of TA-TiTok with 32 tokens. Each latent token attends to prominent semantic and swapping the code leads to appearance changes in the corresponding semantic entity that the latent token focuses on.
Latent Token Attention Maps Reveal Semantic Focus. In the first row of Fig. 4, we overlay the attention maps of each latent token onto the image, where attention is computed from the encoder’s final layer. We observe that each latent token focuses on semantically meaningful regions within the image, rather than specific spatial locations or grid cells as in 2D tokenization. For instance, individual latents capture distinct semantic elements such as leaf, the bird’s head, and the bird’s torso. Notably, we find that at least one latent token (e.g., Latent 11) attends to the entire image, and subsequent experiments reveal that it captures the overall visual style.
Semantic Manipulation Through Latent Code Swapping. Fig. 4 also presents reconstruction results when we swap the code for latent tokens. Interestingly, swapping the code leads to changes in the semantic element corresponding to each latent token, with appearance changes ranging from low-level visual features (e.g., color changes) to high-level semantics (e.g., new objects, object removal, pose changes). For instance, swapping the code for the 4th latent token, which focuses on the leaf, leads to the removal of the leaf or its transformation into a bird or flower. Swapping for the 28th latent token results to the change in bird’s head in terms of its spices or pose, and for 30th latent token swapping results to color change of bird’s torso. Unlike other latent tokens, swapping the 11th latent token focusing on the overall image leads to the change of overall style of the image. For instance, swapping to code 29 leads to the whole scene being redrawn in a flat, vector-graphic style; swapping to code 593 softens high-frequency details; swapping to code 9 produces a purple-tinted stylistic transformation.